OpenStudy (anonymous):

a^5 time B^6 over A^4 time B

OpenStudy (anonymous):

I need help FAST

OpenStudy (anonymous):

are 'a' and 'A' same

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when you are dividing two exponent with same base, you subtract the exponent A^(5-4)*B^6

OpenStudy (anonymous):

A^1 x b^6?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

hold on, I did not see the B

OpenStudy (anonymous):

it should be A^1 b^5

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

\[\frac{A^5B^6}{A^4B}=AB^5\]